1. Solve the differential equation using the Laplace Transform. y + y = V2 sin(V2t), y(0) = 10, y(0) = 0
Answer #1

0 st Given, y+y = y (0) = 10 40)=0 2 Sin (12t), => h{9+y} = 1 {ra sin(83+) sonlar prilago (sy(s) - SYO) - Y(0)) + y (5) siY(s) 1052 +22 (3+2)(2+1) (by poortial fractions using a calculator Y(s) 12. s²+1 2 s²+2 (Hr)DEBE Y(s) = 12 -12 12 3+ (1) 2 3